OpenStudy (anonymous):

an equation of a circle is given. a determine its center. b. determine its radius c. find 2 points on the circle. x^2 + (y-8)^2= 18

OpenStudy (anonymous):

so far, I have a. (0,8) b, approx 4.24 or sq rt 18

OpenStudy (anonymous):

c. insert any real number for x and solve for y or the opposite

OpenStudy (anonymous):

if x^2 > 18, you wont get the point

OpenStudy (anonymous):

@extraven after inserting a real no for x and getting the corresponding real no for y, what we will get, radius or a point on circle?

OpenStudy (anonymous):

we get a point on the circle, the distance between its center and the point you get will be its radius

OpenStudy (anonymous):

@extraven thanks alot. i am also a beginner :) thanks alot

OpenStudy (anonymous):

you welcome :)

OpenStudy (anonymous):

a circle x^2 + y^2 = r^2 with centre at (0,0) and radius r (x-a)^2 + (y-b)^2 = r^2 with centre at (positive a,positive b) and radius r when you see x^2 you can look at it as (x-0)^2

OpenStudy (anonymous):

why do you have to enter a number in for x that is less than 18?

OpenStudy (anonymous):

anything (y-8) ^2 will always come out positive, since, it can either be negative squared or positive squared, both of which will be a positive then you get x^2 + something positive = 18 the positive cannot be negative so the lowest value is 0 then it becomes x^2 + 0 = 18 and you can see why x is limited if the equation is to hold

OpenStudy (anonymous):

ok. When I enter 3=x, I end up with sq rt 9 + 8=y.. can I simplify that to sq rt 17?